Grade 12th passMagical Mathematics[Interesting Approach]

Two balls are drawn at random from an urn containing 5 white and 3 red balls. The first ball is not replaced before the second ball is drawn. Find the probability that the first ball is red, conditional on the second ball being red.
i am getting 2/7. can someone pls clarify

shiya
7 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
Dear student
 
required probability = (3/8 * 2/7) /[(3/8 * 2/7) + (5/8* 3/7)]
 
 = 6/21 = 2/7 is correct answer
 
Regards
Arun